# Fast yet Safe: Early-Exiting with Risk Control

### v1

@@ 1 Introduction
In this work, we address the EENN’s efficiency vs. performance trade-off via statistical frameworks of risk control (RC) [4, 8]. By tuning the EENN’s exiting mechanism based on a user-specified notion of risk, RC aims to enhance the safety of early-exit outputs. We consider several risks that quantify the difference between the early-exit and full model’s outputs, both in terms of prediction quality and uncertainty estimation. Moreover, we study RC frameworks that control the risk with varying degrees of stringency (i.e., in expectation vs. with high probability). We demonstrate the effectiveness of this light-weight, post-hoc solution across a range of tasks, including image classification, semantic segmentation, language modeling, and image generation with diffusion. In particular, we make the following contributions:
@@
  • We improve upon prior work for language modeling [63], demonstrating that our adaptions of risk control allow for less conservative early-exiting, resulting in larger efficiency gains (§ 3.3, § 5.3).
  • We apply, for the first time, early-exiting with risk control to tasks of image classification, semantic segmentation, and image generation (§ 5).
@@ 2 Background
Statistical frameworks for risk control (RC) [1, 4, 8] aim to improve prediction reliability by equipping threshold-based models with safety assurances. Specifically, consider a pre-trained prediction model $\hat{f}{\lambda}$ whose outputs depend on a threshold $\lambda$. For example, given a classification task, the set predictor $\hat{f}{\lambda}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ described by Bates et al. [8] includes a class label in the set if its probability exceeds the threshold, i.e., $\hat{f}_{\lambda}(\bm{x}):=\{\bm{y}\in\mathcal{Y}:\hat{p}(\bm{y}|\bm{x})\geq\lambda\}$. Next, a notion of error for $\hat{f}_{\lambda}$ is captured by defining a problem-specific loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For instance, a meaningful choice for the set predictor could be the miscoverage loss $\ell(\hat{f}_{\lambda}(\bm{x}),\bm{y})=\mathbbm{1}[\bm{y}\notin\hat{f}_{\lambda}(\bm{x})]$, where $\mathbbm{1}[\cdot]$ is the indicator function. The risk associated with a particular threshold $\lambda\in\Lambda$ is then defined as the expected loss
@@ 3.2 Early-Exiting Risks
When calibration labels are present, these can be used measure the early-exit performance through supervised losses. Let $\hat{\bm{o}}_{l}(\bm{x})$ denote a general EENN output for some input $\bm{x}$. It takes the form $\hat{\bm{y}}_{l}$ for predictions and $\hat{p}_{l}(\mathbf{y}|\bm{x})$ for the underlying predictive distribution. The supervised performance gap risk is then defined as
@@ 5 Experiments
We empirically validate early-exiting via risk control on a suite of different tasks: image classification (§ 5.1), semantic segmentation (§ 5.2), language modelling (§ 5.3), and image generation with diffusion (§ 5.4). Our code is publicly available at https://github.com/metodj/RC-EENN. We begin by outlining our general risk control design and evaluation metrics.
@@
We target control of the performance gap and consistency risks defined in § 3.2. For predictions $\hat{\bm{y}}_{l}$ we employ target-specific losses, and, when applicable, for predictive distributions $\hat{p}_{l}(\mathbf{y}|\bm{x})$ our Brier score formulation. We denote these four risks in short as $\mathcal{R}^{G}(\hat{\bm{y}})$, $\mathcal{R}^{G}(\hat{p})$, $\mathcal{R}^{C}(\hat{\bm{y}})$ and $\mathcal{R}^{C}(\hat{p})$. Risk control requirements of different strength are assessed by varying the risk level $\epsilon$. For risk control with high probability, we set $\delta=0.1$ (i.e., 90 % probability). Reported numbers are averaged across multiple trials of calibration and test splitting ($S=100$) to account for sampling effects. Additional results across experiments can be found in Appendix D.
@@ 5.3 Language Modeling
For this task, we replicate the main experiments from Schuster et al. [63] (CALM), using their early-exit version of the T5 model [55] for text summarization on CNN/DM [32] and question answering on SQuAD [56]. Recall that CALM makes use of the Learn-then-Test [1] (LTT) framework for early-exit prediction control, whereas we suggest the Upper Confidence Bound [8] (UCB). In contrast to their experiments which involve excessively large calibration sets ($n\approx 8000$), we explore more practical settings of low calibration sample counts with $n\in\{100,1000\}$. Our results for the performance gap risk (Eq. 6) based on task-specific losses (ROUGE-L for CNN/DM and Token-F1 for SQuAD) are displayed in Fig. 2 and § D.3. In all cases, UCB exit thresholds provide larger computational savings over LTT, while ensuring the same risk control with high probability (Eq. 5). Since particularly pronounced for $n=100$, these results highlight the need for careful framework selection in order to minimize the cost of providing guarantees. Once more, risk control in expectation (CRC, Prop. 1) permits faster exiting due to its weaker safety assurance. Encouragingly, even with as few as $n=100$ calibration samples, CRC exit thresholds reach near-optimal exiting, as indicated by their proximity to the ideal (diagonal) risk line. This suggests that even for modern language tasks, equipping an EENN with notions of safety does not necessitate a (large) compromise on inference efficiency.
@@ 5.4 Image Generation with Early-Exit Diffusion
6 Conclusion, Limitations and Future Work
Our work addresses how to select a ‘safe’ exiting mechanism for early-exit neural networks (EENNs). We propose balancing the EENN’s efficiency vs. performance trade-off via risk control. We empirically validate our light-weight, post-hoc solution on a variety of tasks, and improve upon previous work on early-exit language modeling [63] in several ways. The main limitation of our work is the assumption of a shared exit threshold (Eq. 1). Relaxing this assumption by adopting techniques for high-dimensional thresholds [58, 69] could lead to further efficiency gains. Additionally, multiple risk control extensions provide natural avenues for future work. Firstly, risk control as used in our work is achieved only marginally across observations (Eq. 3), and one could aspire for more granular exit-conditional control [58]. Secondly, the employed risk control frameworks define risk in terms of the expected loss. One could instead aim to control the tails of the loss, e.g., via specific quantiles [65]. Lastly, relaxing the i.i.d assumption on calibration and test data could help extend risk-controlling EENNs to settings with test-time distribution shift [83]. Our work presents a foundation on which to develop such extensions, and is thus an important step towards models that are fast yet safe.
@@ 6 Conclusion, Limitations and Future Work
We thank Mona Schirmer, Rajeev Verma, Christoph-Nikolas Straehle, Patrick Forré and Stephen Bates for helpful discussions and clarifications. This project was generously supported by the Bosch Center for Artificial Intelligence.

### v2

@@ 1 Introduction
In this work, we address the EENN’s efficiency vs. performance trade-off via statistical frameworks of risk control (RC) [4, 8]. By tuning the EENN’s exiting mechanism based on a user-specified notion of risk, RC aims to enhance the safety of early-exit outputs. We consider several risks that quantify the difference between the early-exit and full model’s outputs, both in terms of prediction quality and uncertainty estimation. Moreover, we study RC frameworks that control the risk with varying degrees of stringency (i.e., in expectation vs. with high probability). We demonstrate the effectiveness of this light-weight, post-hoc solution across a range of tasks, including image classification, semantic segmentation, language modeling, image generation with diffusion, and speculative decoding in large language models. In particular, we make the following contributions:
@@
  • We improve upon prior work for language modeling [65], demonstrating that our adaptions of risk control allow for less conservative early-exiting and result in larger efficiency gains (§ 3.3, § 5.3).
  • We apply, for the first time, risk control to early-exiting in image classification, semantic segmentation, and image generation, as well as speculative decoding in large language models (§ 5).
@@ 2 Background
Statistical frameworks for risk control (RC) [1, 4, 8] aim to improve prediction reliability by equipping threshold-based models with safety assurances. Specifically, consider a pre-trained prediction model $\hat{f}_{\lambda}$ whose outputs depend on a threshold $\lambda$. For example, given a classification task, the set predictor $\hat{f}_{\lambda}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ described by Bates et al. [8] includes a class label in the set if its probability exceeds the threshold, i.e., $\hat{f}_{\lambda}(\bm{x}):=\{\bm{y}\in\mathcal{Y}:\hat{p}(\bm{y}|\bm{x})\geq\lambda\}$. Next, a notion of error for $\hat{f}_{\lambda}$ is captured by defining a problem-specific loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For instance, a meaningful choice for the set predictor could be the miscoverage loss $\ell(\hat{f}_{\lambda}(\bm{x}),\bm{y})=\mathbbm{1}[\bm{y}\notin\hat{f}_{\lambda}(\bm{x})]$, where $\mathbbm{1}[\cdot]$ is the indicator function. The risk associated with a particular threshold $\lambda\in\Lambda$ is then defined as the expected loss
@@ 3.2 Early-Exiting Risks
When calibration labels are present, these can be used to measure the early-exit performance through supervised losses. Let $\hat{\bm{o}}_{l}(\bm{x})$ denote a general EENN output for some input $\bm{x}$. It takes the form $\hat{\bm{y}}_{l}$ for predictions and $\hat{p}_{l}(\mathbf{y}|\bm{x})$ for the underlying predictive distribution. The supervised performance gap risk is then defined as
@@ 5 Experiments
We empirically validate early-exiting via risk control on a suite of different tasks: image classification (§ 5.1), semantic segmentation (§ 5.2), language modelling (§ 5.3), image generation with diffusion (§ 5.4), and speculative decoding (§ 5.5). Our code is publicly available at https://github.com/metodj/RC-EENN. We begin by outlining our general risk control design and evaluation metrics.
@@
We target control of the performance gap and consistency risks defined in § 3.2. For predictions $\hat{\bm{y}}_{l}$ we employ target-specific losses, and, when applicable, for predictive distributions $\hat{p}_{l}(\mathbf{y}|\bm{x})$ our Brier score formulation. We denote these four risks in short as $\mathcal{R}^{G}(\hat{\bm{y}})$, $\mathcal{R}^{G}(\hat{p})$, $\mathcal{R}^{C}(\hat{\bm{y}})$ and $\mathcal{R}^{C}(\hat{p})$. Risk control requirements of different strength are assessed by varying the risk level $\epsilon$. Note that our approach is entirely post-hoc, and our experiments leverage existing pretrained EENNs when possible. Thus the underlying models are typically not modified, and we omit the commonly seen performance vs. FLOP curves to instead appropriately benchmark against different risk levels. For risk control with high probability, we set $\delta=0.1$ (i.e., 90 % probability). Reported numbers are averaged across multiple trials of calibration and test splitting ($S=100$) to account for sampling effects. Additional results across experiments can be found in Appendix D.
@@ 5.3 Language Modeling
For this task, we replicate the main experiments from Schuster et al. [65] (CALM), using their early-exit version of the T5 model [57] for text summarization on CNN/DM [33] and question answering on SQuAD [58]. Recall that CALM makes use of the Learn-then-Test [1] (LTT) framework for early-exit prediction control, whereas we suggest the Upper Confidence Bound [8] (UCB). In contrast to their experiments which involve excessively large calibration sets ($n\approx 8000$), we explore more practical settings of low calibration sample counts with $n\in\{100,1000\}$. Our results for the performance gap risk (Eq. 6) based on task-specific losses (ROUGE-L for CNN/DM and Token-F1 for SQuAD) are displayed in Fig. 2 and § D.3. In all cases, UCB exit thresholds provide larger computational savings over LTT, while ensuring the same risk control with high probability (Eq. 5). Since particularly pronounced for $n=100$, these results highlight the need for careful framework selection in order to minimize the cost of providing guarantees. Once more, risk control in expectation (CRC, Prop. 1) permits faster exiting due to its weaker safety assurance. Encouragingly, even with as few as $n=100$ calibration samples, CRC exit thresholds reach near-optimal exiting, as indicated by their proximity to the ideal (diagonal) risk line. This suggests that even for modern language tasks, equipping an EENN with notions of safety does not necessitate a strong compromise on inference efficiency.
@@ 5.4 Image Generation with Early-Exit Diffusion
5.5 Speculative Decoding for Large Language Models
While we primarily focus on early-exiting, our final experiment highlights how risk control can also be applied to other techniques for efficient inference. Here we consider accelerating the inference of large language models (LLMs) using the (soft) speculative decoding approach BiLD [41]. BiLD uses a small draft model to generate multiple tokens autoregressively while the original LLM is only employed for verification.
Figure 6: Empirical test risk (top) and efficiency gains (bottom) for the BiLD model [41] for a machine translation task (De–En) on IWSLT [14] (with $n=500$). We fix the fallback threshold to $0.5$, and apply risk control to the rollback threshold. Our adaptation of UCB again outperforms LTT by yielding larger efficiency gains under the same guarantees. Shading denotes the standard deviation across $S=100$ calibration/test splits.
This step can be performed in a single forward pass for all proposed tokens, necessitating less computations from the larger, more expensive model. During verification the difference in token distributions between the models is computed, and the tokens generated by the draft model are rejected if the difference exceeds a tolerated threshold, triggering a “rollback” (see Eq. 3 in [41]). We apply our risk control frameworks to this rollback threshold, which dictates (i) the similarity of generated text to the output produced solely by the original LLM, and (ii) the associated inference speed-ups in terms of sentences (or samples) per second. Our results in Fig. 6 for the performance gap risk (Eq. 6), as defined via the difference in sentence-level BLEU scores, corroborate our previous findings for language modeling (Fig. 2 and § 5.3). That is, our approaches via CRC (Prop. 1) and UCB (Prop. 2) provide meaningful speed-ups and improve upon the LTT method [1] used by CALM [65], while maintaining the desired risk control across test samples.
6 Discussion
Our work addresses how to select a ‘safe’ exiting mechanism for early-exit neural networks (EENNs). We propose balancing the EENN’s efficiency vs. performance trade-off via risk control, ensuring that accelerated inference does not compromise the quality of the early-exit outputs. We validate our light-weight, post-hoc solution on a variety of tasks and improve upon prior work [65] (see § 4).
Limitations and Future Work.
A key limitation of our work is the reliance on a single shared exit threshold among layers (Eq. 1). While using a shared exit threshold is common [69, 48, 40, 76, 83, 77], relaxing this condition could lead to further efficiency gains. This, however, introduces new challenges both in terms of theory (e.g., defining monotonicity requirements) and practice (e.g., substantially larger search spaces). Overcoming them by adopting recently proposed risk control techniques for high-dimensional thresholds [60, 71] could prove promising for future extensions. Another (simple) workaround is to reduce the multi-dimensional problem back to a single threshold by use of a threshold function, as partially explored in [65]. Instead of working directly with multiple thresholds, our risk control framework can then be applied to this scalar parameter. Additionally, multiple risk control extensions provide natural avenues for future work. Firstly, risk control as used in our work is achieved only marginally across observations (Eq. 3), and one could aspire for more granular exit-conditional control [60]. Secondly, the employed risk control frameworks define risk in terms of the expected loss. One could instead aim to control the tails of the loss, e.g., via specific quantiles of interest [67]. Lastly, relaxing the i.i.d assumption on calibration and test data could help extend risk-controlling EENNs to scenarios with test-time distribution shifts [86] or to online updating strategies [23].
Broader Impacts.
EENNs provide a simple and effective approach to dynamic computation [29]. Their computational savings can reduce energy costs and the carbon footprint, as well as allow the model to be deployed on resource-constrained hardware. By incorporating a ’safe’ exit mechanism into these models, we improve their trustworthiness and strengthen the reliability of their intermediate outputs, along with any decisions based on them. This facilitates safer model deployment in real-world applications and contributes to more responsible decision-making. While we do not foresee any direct negative consequences from our work, improper use of our risk control framework can lead to violations or misinterpretations of its provided guarantees. This, in turn, can risk instilling a false sense of security. Overall, we believe that our work outlines an effective approach to improve the reliability of EENNs and to safely balance their inherent efficiency vs. performance trade-off. In doing so, it contributes to the goal of developing models that are ultimately fast yet safe.
@@ 6 Discussion
We thank Mona Schirmer, Rajeev Verma, Christoph-Nikolas Straehle, Patrick Forré and Stephen Bates for helpful discussions and clarifications. We are also grateful to the anonymous reviewers who helped improve the work with their constructive feedback. This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not utilize resources from Johns Hopkins University for this project.
@@ D.3 Language Modeling
Figure 12: Empirical test risk (top) and efficiency gains (bottom) for the CALM model for text summarization on CNN/DM across different confidence measures (see Schuster et al. [65], §3.5). From left to right: Hidden state saturation, meta-classifiers, and top-2 softmax difference. Our employed risk control frameworks based on CRC and UCB continue to outperform LTT across all measures of confidence. Shading denotes the standard deviation across $S=100$ splits.